# Differential phase contrast from electrons that cause inner shell ionization Michael Deimetry Timothy C. Petersen

### v1

@@ Abstract
Differential Phase Contrast (DPC) imaging, in which deviations in the bright field beam are in proportion to the electric field, has been extensively studied in the context of pure elastic scattering. Here we discuss differential phase contrast formed from core-loss scattered electrons, i.e. those that have caused inner shell ionization of atoms in the specimen, using a transition potential approach for which we study the convergence properties. In the phase object approximation, we show formally that this is mainly a result of preservation of elastic contrast. Through simulation we demonstrate that whether the inelastic DPC images show element selective contrast depends on the spatial range of the ionization interaction, and specifically that when the energy loss is low the delocalisation can lead to contributions to the contrast from atoms other than that ionized. We further show that inelastic DPC images remain robustly interpretable to larger thicknesses than is the case for elastic DPC images, owing to the incoherence of the inelastic wavefields, though subtleties due to channelling remain.
@@ 1 Introduction
This paper explores the imaging formation dynamics of DPC imaging from core-loss-filtered 4D STEM, describing core-loss scattering via the transition potential formulation [21, 23, 25, 26]. Our exploration is centred around two key results. The first is the experimental finding of Haas and Koch [20] that the zero-loss-filtered and core-loss-filtered DPC images have qualitatively similar appearance, as reproduced in the simulated DPC images for hBN in Fig.1(b) and (c) respectively. In particular, the nitrogen atoms are clearly visible in the core-loss-filtered DPC image for the boron K-edge, and so, in this case, core-loss-filtered DPC imaging is not providing element-specific contrast. The second, demonstrated in the simulated DPC images for 200 Å thick SrTiO3 in Fig.1(d) and (e), is that the image contrast in core-loss-filtered DPC imaging (Fig.1(e)) is more robust with respect to thickness (and defocus) than that in zero-loss-filtered DPC imaging (Fig.1(d)).
@@
Figure 1: (a) Simplified schematic of the energy-filtered 4D STEM setup. An electron probe is raster scanned across a monolayer of hBN. The electrons pass through a spectrometer with a slit chosen to only let through electrons with kinetic energies in a desired energy range.
(b) Elastic and (c) inelastic (with energy filter 50 eV above the boron K-shell edge) simulated DPC images for monolayer hBN. The nitrogen columns are more pronounced in the elastic image, but still clearly visible in the inelastic image from the boron K-shell. (d) Elastic and (e) inelastic (with energy filter 50 eV above the Ti -shell) DPC images for 200 Å thick . Whereas at this thickness it is difficult to interpret the structure clearly from the elastic image, the inelastic image is more robustly interpretable, with the Ti column locations showing clear differential phase contrast. These simulations assume 300 keV beam electrons and a 15.7 mrad probe-forming convergence semiangle.
@@ 2 Transition potential formulation of core-loss scattering
Figure 2: Schematic of the inelastic scattering process in a monolayer of atoms. The incident wavefield produces an inelastic wave with probability governed by the transition potential . This procedure is repeated for each final state of each atom of the ionized species, and the resultant diffraction patterns are summed incoherently.
@@
where are the real-space coordinates transverse to the optic axis, (in which is the Lorentz factor, is the electron rest mass, is the relativistically-corrected de-Broglie wavelength and is the reduced Planck’s constant), and is a (projected) transition potential222In the literature is also referred to as a projected matrix element, Mller potential or simply a matrix element. The choice of units also varies, with other choices leading to a prefactor that differs accordingly from the of Eq.1.. Since for the single plane of atoms depicted in Fig.2 the wavefield is at the exit surface of the sample, its contribution to the diffraction pattern is given by the intensity of its Fourier transform. The principles of quantum mechanics oblige us to (incoherently) sum all contributions from sample final states that are not directly observed but which are consistent with what is measured. Thus the core-loss-filtered diffraction intensity may be written
@@
where is the elementary charge and is the permittivity of free space. Equation3 corresponds to the coupling of the bound state and the continuum state of the atomic electron ejected during the ionization event.333We will not consider bound-bound transitions, though they can be described by a similar formalism [28]. The projected transition potential appearing in Eq.1 is then obtained by [21, 25]
@@
Writing the bound and continuum states explicitly as a product of angular and radial terms444Following Ref. [29], we use relativistic Hartree-Fock wave functions for the bound states and Hartree-Slater wave functions for the continuum states., and , Eq.3 can be simplified to
@@
(in which are spherical Bessel functions of order of the first kind) can be computed efficiently using the fast Hankel transform555The Hankel transform is defined as
@@
Figure3 shows various titanium L2,3-shell excitation probabilities. The greatest amplitude is near the atom, implying beam electrons closest to the titanium atom are more likely to cause ionization events. The vortex phase structure occurs because, despite the sum over in Eq.6, all the azimuthal dependence is contained in a phase factor of the form . The so-called winding number of the vortex is given by , where positive values imply increasing phase in the clockwise direction. The transitions are related by the magnetic quantum numbers through the relationship
@@
Figure 3: L2,3-shell titanium transition probabilities, , for beam energy 80keV and eV. The magnitude, depicted as brightness, shows that beam electrons near the atom are most likely to cause ionization events. The phase, depicted modulo by colour, has a vortex structure with the winding number given by .
The transition potentials in each tile in Fig.3 are normalised to a common maximum magnitude to make the phase structure clear. However, this obscures that the magnitudes vary significantly. In particular, they tend to decrease with increasing . Transitions which are unlikely to be realized can be neglected with little consequence on the accuracy of the calculation. Indeed, simulation relies on being able to reduce the in-principle infinite number of final states to sum over in Eq.2 ( being any non-negative integer) to a limited number that capture the majority of the inelastic scattering probability. The probability of inelastic scattering into different final states is, for the same incident wavefield, bounded666This follows by applying the triangle inequality to the integrated intensity on both sides of Eq.1, an idea developed further in the Supplementary Material. by the integrated intensity in . We therefore consider the rate of convergence of the following sequence in :
A remarkable pattern emerges when the sequence (normalised by the term) is plotted on a log scale as illustrated in Fig.4. Beginning from for K-shell and for -shell, the terms of the sequence fall off exponentially (linear behaviour with the vertical scale plotted logarithmically). This fall-off varies depending on several experimental parameters: the atomic number is most significant, followed by the energy loss above ionization threshold . The dependence on accelerating voltage is not shown because it is very weak indeed.
As derived in the supplementary material, an approximate analytic expression for the fall-off observed for the K-shell is given by
@@
Figure 4: Sequence in Eq.10 (normalised by the term) plotted for boron and zinc for ionization energies 50 eV and 200 eV above the (a) K- and (b) - shell edge. Solid lines are a guide to the eye, emphasising the linear trend. Dotted lines show the prediction of the approximate analytic expression, Eq.11, for K-shell ionization.
@@ 3 Case study – monolayer materials
This inelastic DPC expression is very similar in form to the elastic CoM expression in Eq.15, save now that the phase gradient is modulated by the inelastic transition probability. In particular, the contrast does not arise from the inelastic scattering (interpreted to be the transition potential) but rather from the elastic scattering (interpreted to be , the phase of the elastic scattering transmission function). The result can thus be considered an example of preservation of elastic contrast [32, 33], where contrast deriving from the elastic scattering is retained in the inelastic signal.
Strictly speaking, Eq.19 means the DPC image is no longer an exact gradient. However, if the inelastic transition probability is approximately constant in the vicinity of the atom sites then
@@
The interpretation of Eqs.19 and 21 depends on the spatial extent of the transition probabilities . If they are so delocalised that would be approximately constant throughout the material then the inelastic DPC signal would be proportional to the elastic DPC signal, with contrast from all atoms visible. (This is consistent with the results of Beyer et al. [35] for plasmon-loss-filtered, atomic-resolution DPC.) If they are sufficiently localised as to become negligible over the interatomic distance in the direction transverse to the beam axis then only the columns containing the ionized elements will appear in the image, producing element selective contrast. If they are between these limits, being non-zero but reduced in magnitude at adjacent atomic sites, then those sites will appear in the EFDPC image with reduced contrast (relative to the elastic case).
@@
The spatial extent of the transition probabilities depends on accelerating voltage of the beam, the ionization edge (i.e. the atomic species and the initial state), and the energy loss above the atom’s ionization threshold, . This is explored in the EFDPC image tableaus over and atomic number in Fig.5(a) and (b). In the case Z=5 the structure is monolayer hBN; for other Z we assume a fictitious structure in which the boron atoms are replaced by atoms of that atomic number. In Fig.5(a), the K-shell ionization of that atom is assumed; in Fig.5(b), L1-shell ionization is assumed. The transition potentials become more localized with increasing atomic number and/or energy above threshold, and consequently the contrast at the nitrogen atom sites, indicated by arrows, become fainter.
@@
Hence, for element-selective contrast corresponding to a low atomic number () the energy window needs to be several tens of electron volts above the ionization edge. For larger atomic numbers, the energy window can be moved closer to the ionization edge. Since fewer inelastic events will occur further from the ionization edge, lower atomic numbers present the greatest challenge experimentally. The experimental images of Haas and Koch [20] of boron K-shell ionization do not show boron-selective contrast, instead having contrast evident at the nitrogen sites too. This is expected since for element-selective contrast for boron the energy window need to be placed around 50 eV above the K-edge. However, since this distance from the edge will yield much lower signal-to-noise, the energy filter for that experiment was likely positioned a few electron volts beyond the boron K-edge.
@@
Figure 5: (a) K-shell and (b) -shell EFDPC tableaus assuming a hBN-like structure and varying , the atomic number of the atoms at what in hBN would be the boron site, and . These simulations assume 300 keV beam electrons and a 15.7 mrad probe-forming convergence semiangle. The arrows indicate nitrogen sites where there is a gradual reduction in contrast with increasing (of the atom being ionized) and/or energy above threshold. To convey in more detail the spatial localization of the transition potentials, we show what percentage of the transition probability (integrated mod-square of the transition potential) is contained within a given radius (see Eq.22) for (c) K-shell (d) -shell for all possible combinations of , eV and beam energy keV up to inclusive.
@@
for for various . Since the final CoM image will contain contributions from all allowed intermediate states, Eq.22 does not distinguish the individual transition contributions to the overall fraction of all expected transitions, hence the summation over final states. Figure5(c) and (d) contain the solution of Eq.22 for all combinations of the accelerating voltage keV, the ionization energy above threshold eV, the atomic number and the K, L1 shells. The summation in Eq.22 is truncated at inclusive since, as implied by Fig.4, the remaining transitions are heavily suppressed. For small atomic numbers both beam energy and are significant in the potential’s localization albeit is dominant. However, this dependence gradually flips for larger atomic numbers. The same is true, but more pronounced, for higher shells.
@@ 4 Case study – thick specimen
Figure 6: Schematic of the inelastic scattering process in a thick specimen of condensed matter. The incident wavefield is elastically propagated using the scattering matrix to the depth of the atom being ionized, produces an inelastic wave with probability governed by the transition potential , and is then elastically propagated using the scattering matrix to the exit surface. This procedure is repeated for each final state of each atom of the ionized species, and the resultant diffraction patterns are summed incoherently.
@@
Figure 7: Thickness-defocus tableaus of elastic DPC images (left) and EFDPC 10 eV beyond the Ti L1-edge (right), where underfocus is negative. All images assume a 15.7 mrad probe-forming aperture and 300 keV beam energy. With elastic DPC normalized by thickness and EFDPC by the thickness squared here, the contrast within each unit cell should be independent of thickness within the phase-object approximation’s domain of validity. This contrast is seen to decrease beyond about 35 Å for both regimes, evidence of the phase-object approximation starting to break down.
@@
In the inelastic case, Eq.19, in addition to the linear dependence of on sample thickness, the incoherent summation of contributions from the atoms along the column is also linear in thickness. Hence, normalising the EFDPC images by the square of the sample thickness will lead to the contrast within each unit cell being independent of thickness within the phase object approximation’s domain of applicability. This is seen in the EFDPC tableau for the Ti L1-edge in the right panel in Fig.7. The quantitative domain of validity of the phase object approximation is seen to be consistent with the elastic case, with very similar contrast seen for thicknesses below 35 Å but a decrease in contrast at larger thicknesses. However, excepting for large defocus in the thinnest sample, the pattern in the EFDPC image persists to larger thicknesses than in the elastic case.
@@
To show that this qualitative robustness persists to still larger thicknesses, Fig.8 shows thickness-defocus tableaus for (a) elastic DPC, and for EFDPC of (b) the O-K edge, and (c) Ti L1-edge. These extended tableaus are not thickness-normalised and more clearly show that a defocus value near the specimen midplane tends to maximise the contrast. In contrast to the elastic case, where we see that the defocus-thickness combinations over which the images show clear differential phase contrast is somewhat limited, in the inelastic case the images show clear differential phase contrast over a wide range of thickness and defocus values. Atom-selective contrast is evident, with only the Ti sites clearly visible in the Ti L1-edge case. In the O K-edge case, we see both the pure O and TiO columns, with the latter showing higher contrast due to the larger gradient of the elastic potential on that more-strongly-scattering TiO column. Inelastic DPC thus constitutes an interesting compromise: it contains phase-contrast-like elements through preservation of elastic contrast, but maintains interpretable contrast over a wide thickness as is more usually found in incoherent imaging (somewhat reminiscent of energy filtered TEM [33]).
@@
Figure 8: Thickness-defocus tableaus of for (a) elastic DPC images, (b) EFDPC images 10 eV beyond the O K-edge, and (c) EFDPC images 10 eV beyond the Ti L1-edge, assuming a 15.7 mrad probe-forming aperture and 300 keV beam energy, and where underfocus is negative. Elastic DPC remains interpretable for specimens of thickness of around 40 Å given the probe is focused into the mid-plane of the specimen. EFDPC remains interpretable throughout the entire range observed. The emphasised tiles in (c) correspond to the parameters used in Fig.9.
@@
However, this robust interpretability does not mean that channelling effects are absent. To see this, we decompose the incoherent sum over final states in Eq.17 to the total contribution from unit-cell-thick layers perpendicular to the optic axis. Figure9(a) shows this layer decomposition for the tiles in the coloured boxes in Fig.8(c), these defocus-thickness parameter combinations — , , and — being diverse choices over the parameter space shown. We see that the contribution from different layers varies. The layers which show the clearest differential phase contrast are also those which contribute most to the signal, which helps explain the robustness of the total contrast, even though the contribution from some layers is far less directly interpretable. Which layers contribute the most varies widely for the different defocus-thickness parameter combinations: in Fig.8(b) the layers near the midplane dominate the intensity; in Fig.8(c) the layers prior to the midplane dominate; in Fig.8(d) the layers prior and subsequent to the mid-plane provide comparable contributions. Note too that the layers which contribute the most are not necessarily those at the nominal in-focus plane. Nevertheless, in each case the sum over layers produces a fairly similar total image, consistent with the fairly uniform contrast in Fig.8(c).

### v2

@@ Abstract
Differential Phase Contrast (DPC) imaging, in which deviations in the bright field beam are in proportion to the electric field, has been extensively studied in the context of pure elastic scattering. Here we discuss differential phase contrast formed from core-loss scattered electrons, i.e. those that have caused inner shell ionization of atoms in the specimen, using a transition potential approach for which we study the number of final states needed for a converged calculation. In the phase object approximation, we show formally that differential phase contrast formed from core-loss scattered electrons is mainly a result of preservation of elastic contrast. Through simulation we demonstrate that whether the inelastic DPC images show element selective contrast depends on the spatial range of the ionization interaction, and specifically that when the energy loss is low the delocalisation can lead to contributions to the contrast from atoms other than that ionized. We further show that inelastic DPC images remain robustly interpretable to larger thicknesses than is the case for elastic DPC images, owing to the incoherence of the inelastic wavefields, though subtleties due to channelling remain. Lastly, we show that while a very high dose will be needed for sufficient counting statistics to discern differential phase contrast from core-loss scattered electrons, there is some enhancement of signal-to-noise ratio with thickness that makes inelastic DPC imaging more achievable for thicker samples.
@@ 1 Introduction
This paper explores the imaging formation dynamics of DPC imaging from core-loss-filtered 4D STEM,222All DPC images shown are calculated by evaluating the first moment assuming a pixel detector, though we expect results from a segmented detector to be similar, as is the case for elastic DPC imaging [25, 26]. describing core-loss scattering via the transition potential formulation [21, 23, 27, 28]. Our exploration is centred around two key results. The first is the experimental finding of Haas and Koch [20] that the zero-loss-filtered and core-loss-filtered DPC images have qualitatively similar appearance, as reproduced in the simulated DPC images for hBN in Fig.1(b) and (c) respectively. In particular, the nitrogen atoms are clearly visible in the core-loss-filtered DPC image for the boron K edge, and so, in this case, core-loss-filtered DPC imaging is not providing element-specific contrast. The second, demonstrated in the simulated DPC images for 200 Å thick SrTiO3 in Fig.1(d) and (e), is that the image contrast in core-loss-filtered DPC imaging (Fig.1(e)) is more robust with respect to thickness (and defocus) than that in zero-loss-filtered DPC imaging (Fig.1(d)). While Fig.1(b)-(e) assume infinite dose to emphasise the similarity in appearance between elastic and core-loss-filtered DPC images, in practice core-loss-filtered DPC images will typically have contrast several orders of magnitude smaller than that for elastic DPC images. We will use infinite dose calculations to explore contrast formation mechanisms and imaging dynamics, but return to practical considerations about dose later in the paper.
@@
Figure 1: (a) Simplified schematic of the energy-filtered 4D STEM setup. An electron probe is raster scanned across a monolayer of hBN. The electrons pass through a spectrometer with a slit chosen to only let through electrons with kinetic energies in a desired energy range. (b) Elastic and (c) inelastic (with energy filter 50 eV above the boron K edge) simulated DPC images of the first moment along the -direction (axes inset in (b)), for a detector spanning around 100 mrad in the - and -directions for monolayer hBN. The nitrogen columns are more pronounced in the elastic image, but still clearly visible in the inelastic image from the boron K shell. (d) Elastic and (e) inelastic (with energy filter 50 eV above the titanium L1 edge) DPC images of the first moment along the -direction for a detector with angular range extending out to 35 mrad for 200 Å thick . Whereas at this thickness it is difficult to interpret the structure clearly from the elastic image, the inelastic image is more robustly interpretable, with the Ti column locations showing clear differential phase contrast. These simulations assume 300 keV beam electrons and a 15.7 mrad probe-forming convergence semiangle. As the units for our EFDPC calculations are not intuitive, colour scales for EFDPC images simply show the zero value and distinguish positive from negative (see section 5 for a discussion).
@@ 2 Transition potential formulation of core-loss scattering
Figure 2: Schematic of the inelastic scattering process in a monolayer of atoms. The incident wavefield produces an inelastic wave governed by the transition potential . This procedure is repeated for each final state of each atom of the ionized species, and the resultant diffraction patterns are summed incoherently.
@@
where are the real-space coordinates transverse to the optic axis, (in which is the Lorentz factor, is the electron rest mass, is the relativistically-corrected de-Broglie wavelength and is the reduced Planck’s constant), and is a (projected) transition potential333In the literature is also referred to as a projected matrix element, Mller potential or simply a matrix element. The choice of units also varies, with other choices leading to a prefactor that differs accordingly from the of Eq.1.. Since for the single plane of atoms depicted in Fig.2 the wavefield is at the exit surface of the sample, its contribution to the diffraction pattern is given by the intensity of its Fourier transform. The principles of quantum mechanics oblige us to (incoherently) sum all contributions from sample final states that are not directly observed but which are consistent with what is measured. Thus the core-loss-filtered diffraction intensity may be written
@@
where is the elementary charge and is the permittivity of free space. Equation3 corresponds to the coupling of the bound state and the continuum state of the atomic electron ejected during the ionization event.444We will not consider bound-bound transitions, though they can be described by a similar formalism [30]. The projected transition potential appearing in Eq.1 is then obtained by [21, 27]
@@
Writing the bound and continuum states explicitly as a product of angular and radial terms555Following Ref. [31], we use relativistic Hartree-Fock wave functions for the bound states and Hartree-Slater wave functions for the continuum states., and , Eq.3 can be simplified to
@@
(in which are spherical Bessel functions of order of the first kind) can be computed efficiently using the fast Hankel transform666The Hankel transform is defined as
@@
Figure3 shows various titanium L2,3 shell transition potentials. The greatest amplitude is near the atom, implying beam electrons closest to the titanium atom are more likely to cause ionization events. The vortex phase structure occurs because, despite the sum over in Eq.6, all the azimuthal dependence is contained in a phase factor of the form . The so-called winding number of the vortex is given by , where positive values imply increasing phase in the clockwise direction. The transitions are related by the magnetic quantum numbers through the relationship
@@
Figure 3: L2,3 shell titanium transition potentials, , for beam energy 80keV and eV. The magnitude, depicted as brightness, shows that beam electrons near the atom are most likely to cause ionization events. The phase, depicted modulo by colour, has a vortex structure with the winding number given by .
The transition potentials in each tile in Fig.3 are normalised to a common maximum magnitude to make the phase structure clear. However, this obscures that the magnitudes vary significantly. In particular, they tend to decrease with increasing . Transitions which are unlikely to be realized can be neglected with little consequence on the accuracy of the calculation. Indeed, simulation relies on being able to reduce the in-principle infinite number of final states to sum over in Eq.2 ( being any non-negative integer) to a limited number that capture the majority of the inelastic scattering. The
transition to different final states is, for the same incident wavefield, bounded777This follows by applying the triangle inequality to the integrated intensity on both sides of Eq.1, an idea developed further in the Supplementary Material. by the integrated intensity in . We therefore consider the rate of convergence of the following sequence in :
A remarkable pattern emerges when the sequence (normalised by the term) is plotted on a log scale as illustrated in Fig.4. Beginning from for K shell and for shell, the terms of the sequence fall off exponentially (linear behaviour with the vertical scale plotted logarithmically). This fall-off varies depending on several experimental parameters: the atomic number is most significant, followed by the energy loss above ionization threshold . The dependence on accelerating voltage is not shown because it is very weak, having largely cancelled out in the normalisation by the term. As shown in the Supplementary Material, the dominant energy dependence of the inelastic transition potentials comes through the approximate scaling with : decreasing the beam energy would decrease and thus increase .
As derived in the Supplementary Material, an approximate analytic expression for the fall-off observed for the K shell is given by
@@
Figure 4: Sequence in Eq.10 (normalised by the term) plotted for boron and zinc for ionization energies 50 eV and 200 eV above the (a) K and (b) edge. Solid lines are a guide to the eye, emphasising the linear trend. Dotted lines show the prediction of the approximate analytic expression, Eq.11, for K shell ionization.
@@ 3 Case study – monolayer materials
This inelastic DPC expression is very similar in form to the elastic CoM expression in Eq.15, save now that the phase gradient is modulated by the magnitude squared of the inelastic transition potentials. In particular, the contrast does not arise from the inelastic scattering (interpreted to be the transition potential) but rather from the elastic scattering (interpreted to be , the phase of the elastic scattering transmission function). The result can thus be considered an example of preservation of elastic contrast [34, 35], where contrast deriving from the elastic scattering is retained in the inelastic signal.
Strictly speaking, Eq.19 means the DPC image is no longer an exact gradient. However, if the inelastic transition potential is approximately constant in the vicinity of the atom sites then
@@
The interpretation of Eqs.19 and 21 depends on the spatial extent of the magnitude squared of the transition potentials, . If they are so delocalised that would be approximately constant throughout the material then the inelastic DPC signal would be proportional to the elastic DPC signal, with contrast from all atoms visible. (This is consistent with the results of Beyer et al. [37] for plasmon-loss-filtered, atomic-resolution DPC.) If they are sufficiently localised as to become negligible over the interatomic distance in the direction transverse to the beam axis then only the columns containing the ionized elements will appear in the image, producing element selective contrast.888Eq. (20) would only give an element specific exact derivative if was a top-hat function encompassing all of the elastic potential of the atom of interest and none of the elastic potential of neighbouring atoms. This is not the case, and so Eqs. (20) and (21) are only approximations. Despite this, Figs. 5(a) and (b) clearly show contrast qualitatively consistent with differential phase contrast. This is somewhat reminiscent of the distinction between first moment imaging using a pixel detector and differential phase contrast using a segmented detector: the latter is only an approximation to the former, but suffices for qualitative, and often for semi-quantitative, analysis. If they are between these limits, being non-zero but reduced in magnitude at adjacent atomic sites, then those sites will appear in the EFDPC image with reduced contrast (relative to the elastic case).
@@
The spatial extent of depends on accelerating voltage of the beam, the ionization edge (i.e. the atomic species and the initial state), and the energy loss above the atom’s ionization threshold, . This is explored in the EFDPC image tableaus over and atomic number in Fig.5(a) and (b). In the case Z=5 the structure is monolayer hBN; for other Z we assume a fictitious structure in which the boron atoms are replaced by atoms of that atomic number. In Fig.5(a), the K shell ionization of that atom is assumed; in Fig.5(b), L1 shell ionization is assumed. The transition potentials become more localized with increasing atomic number and/or energy above threshold (though for the lowest energy transition shown, Be L1, the delocalisation remains significant even at eV), and consequently the contrast at the nitrogen atom sites, indicated by arrows, become fainter.
@@
Hence, for element-selective contrast corresponding to a low atomic number () the energy window needs to be several tens of electron volts above the ionization edge. For larger atomic numbers, the energy window can be moved closer to the ionization edge. Since fewer inelastic events will occur further from the ionization edge, lower atomic numbers present the greatest challenge experimentally. The experimental images of Haas and Koch [20] of boron K shell ionization do not show boron-selective contrast, instead having contrast evident at the nitrogen sites too. This is expected since for element-selective contrast for boron the energy window need to be placed around 50 eV above the K edge. However, since this distance from the edge will yield much lower signal-to-noise, the energy filter for that experiment was likely positioned a few electron volts beyond the boron K edge.
@@
Figure 5: (a) K shell and (b) shell EFDPC of the first moment along the -direction for a detector spanning around 100 mrad in the - and -directions tableaus assuming a hBN-like structure and varying , the atomic number of the atoms at what in hBN would be the boron site, and . These simulations assume 300 keV beam electrons and a 15.7 mrad probe-forming convergence semiangle. The arrows indicate nitrogen sites where there is a gradual reduction in contrast with increasing (of the atom being ionized) and/or energy above threshold. Each tile in the tableaus is displayed with the full colour scale spanning its individual minimum to maximum values. To convey in more detail the spatial localization of the transition potentials, we show what percentage of the integrated magnitude squared of the transition potential is contained within a given radius (see Eq.22) for (c) K shell (d) shell for all possible combinations of , eV and beam energy keV up to (a,c) and (b,d) inclusive.
@@
for for various . Since the final CoM image will contain contributions from all allowed intermediate states, Eq.22 does not distinguish the individual transition contributions to the overall fraction of all expected transitions, hence the summation over final states. Figure5(c) and (d) contain the solution of Eq.22 for all combinations of the accelerating voltage keV, the ionization energy above threshold eV, the atomic number and the K, L1 shells. The summation in Eq.22 is truncated at inclusive for Fig.5(c) and inclusive for Fig.5(d) since, as implied by Fig.4, the remaining transitions are heavily suppressed. For small atomic numbers both beam energy and are significant in the potential’s localization albeit is dominant. However, this dependence gradually flips for larger atomic numbers. The same is true, but more pronounced, for higher shells.
@@ 4 Case study – thick specimen
Figure 6: Schematic of the inelastic scattering process in a thick specimen of condensed matter. The incident wavefield is elastically propagated using the scattering matrix to the depth of the atom being ionized, produces an inelastic wave depending on the transition potential , and is then elastically propagated using the scattering matrix to the exit surface. This procedure is repeated for each final state of each atom of the ionized species, and the resultant diffraction patterns are summed incoherently.
@@
Figure 7: Thickness-defocus tableaus of elastic first moment (-direction) DPC images (left) and first moment (-direction) EFDPC 10 eV beyond the titanium L1 edge (right), where underfocus is negative. All images assume a 15.7 mrad probe-forming aperture, a detector extending out to 35 mrad scattering angle and 300 keV beam energy. Within each tableau, all tiles are displayed on the same colour scale such that relative contrast can be compared between tiles. With elastic DPC normalized by thickness and EFDPC by the thickness squared here, the contrast within each unit cell should be independent of thickness within the phase-object approximation’s domain of validity. This contrast is seen to decrease beyond about 35 Å for both regimes, evidence of the phase-object approximation starting to break down.
@@
In the inelastic case, Eq.19, in addition to the linear dependence of on sample thickness, the incoherent summation of contributions from the atoms along the column is also linear in thickness. Hence, normalising the EFDPC images by the square of the sample thickness will lead to the contrast within each unit cell being independent of thickness within the phase object approximation’s domain of applicability. This is seen in the EFDPC tableau for the titanium L1 edge in the right panel in Fig.7. The quantitative domain of validity of the phase object approximation is seen to be consistent with the elastic case, with very similar contrast seen for thicknesses below 35 Å but a decrease in contrast at larger thicknesses. However, excepting for large defocus in the thinnest sample, the pattern in the EFDPC image persists to larger thicknesses than in the elastic case.
@@
To show that this qualitative robustness persists to still larger thicknesses, Fig.8 shows thickness-defocus tableaus for (a) elastic DPC, and for EFDPC of (b) the oxygen K edge, and (c) titanium L1 edge. These extended tableaus are not thickness-normalised and more clearly show that a defocus value near the specimen midplane tends to maximise the contrast. In contrast to the elastic case, where we see that the defocus-thickness combinations over which the images show clear differential phase contrast is somewhat limited, in the inelastic case the images show clear differential phase contrast over a wide range of thickness and defocus values. Atom-selective contrast is evident, with only the Ti sites clearly visible in the titanium L1 edge case. In the oxygen K edge case, we see both the pure O and TiO columns, with the latter showing higher contrast due to the larger gradient of the elastic potential on that more-strongly-scattering TiO column. Inelastic DPC thus constitutes an interesting compromise: it contains phase-contrast-like elements through preservation of elastic contrast, but maintains interpretable contrast over a wide thickness as is more usually found in incoherent imaging (somewhat reminiscent of energy filtered TEM [35]).
@@
Figure 8: Thickness-defocus tableaus of for (a) elastic first moment (-direction) DPC images, (b) first moment (-direction) EFDPC images 10 eV beyond the oxygen K edge, and (c) first moment (-direction) EFDPC images 10 eV beyond the titanium L1 edge. All images assume a detector extending out to 35 mrad scattering angle, a 15.7 mrad probe-forming aperture and 300 keV beam energy, where underfocus is negative. Within each tableau, all tiles are displayed on the same colour scale such that relative contrast can be compared between tiles. Elastic DPC remains interpretable for specimens of thickness of around 40 Å given the probe is focused into the mid-plane of the specimen. EFDPC remains interpretable throughout the entire range observed. The emphasised tiles in (c) correspond to the parameters used in Fig.9.
@@
However, this robust interpretability does not mean that channelling effects are absent. To see this, we decompose the incoherent sum over final states in Eq.17 to the total contribution from unit-cell-thick layers perpendicular to the optic axis. Figure9(a) shows this layer decomposition for the tiles in the coloured boxes in Fig.8(c), these defocus-thickness parameter combinations — , , and — being diverse choices over the parameter space shown. We see that the contribution from different layers varies. The layers which show the clearest differential phase contrast are also those which contribute most to the signal, which helps explain the robustness of the total contrast, even though the contribution from some layers is far less directly interpretable. Which layers contribute the most varies widely for the different defocus-thickness parameter combinations: in Fig.9(b) the layers near the midplane dominate the intensity; in Fig.9(c) the layers prior and subsequent to the mid-plane provide comparable contributions; in Fig.9(d) the layers prior to the midplane dominate. Note too that the layers which contribute the most are not necessarily those at the nominal in-focus plane. Nevertheless, in each case the sum over layers produces a fairly similar total image, consistent with the fairly uniform contrast in Fig.8(c).
Figure 9: (a) Layer contributions (from a sub-set of layers) to the first moment (-direction) EFDPC images for the three different thickness and defocus parameter combinations indicated by the coloured boxes in Fig.8(c), specifically 78.1 Å (20 layers) and surface defocus (top), 156.2 Å (40 layers) and Å defocus (middle), and 234.3 Å (60 layers) and surface defocus (bottom), all assuming specimen, 300 keV beam electrons, 15.7 mrad aperture and energy filter placed 10 eV beyond the Ti edge. A layer is taken to be the depth of a unitcell i.e. 3.9 Å. For each of the three cases respectively, (b), (c) and (d) plot the energy-filtered, integrated DPC (EFiDPC), integrated inelastic intensity (total EELS), scattering cross-section (, Eq.25), and the scattering cross-section with a top-hat function ( w/ top-hat, Eq.24) for the probe on the TiO column as a function of layer depth. Insets show the 2D EFiDPC maps about the TiO column. All plots are scaled such that maxima and minima align, so while the EFiDPC signal given on the left axis and the cross-section given on the right correctly conveys the zero location and the contrast, the scales are in arbitrary units. The broad similarity between the plots results from the underlying channelling of the probe being common to each signal.